# On the Importance of Text Preprocessing for Multimodal Representation Learning and Pathology Report Generation

### v2

@@ Abstract
Keywords:
Vision Language Modeling Histopathology Text Preprocessing
@@ 1 Introduction
Vision-language modeling has seen much improvement in recent years [9, 10, 18, 23]. Following the success in the domain of natural images, similar models have been developed for medical domains such as pathology [1, 5, 13, 19]. Applications of vision-language models in pathology include uni- and cross-modal retrieval of cases from databases and automated report generation. The latter can potentially alleviate the increasing workload of pathologists [2, 8].
@@
In addition to descriptions of cell and tissue appearances on hematoxylin and eosin (H&E)-stained whole slide images (WSIs), pathology reports often also include clinical information, patient history, and additional diagnostic results from immunohistochemical stains and molecular tests. These types of information are either difficult to predict correctly or cannot accurately be inferred from the H&E-stained WSIs at all. If sentences with this information are part of the training dataset, then vision-language models are prone to hallucination (i.e., the generation of statements that contradict or cannot be verified from the source content) [7]. For example, outcomes of molecular testing are present in the reports generated by TITAN [5] in combination with PathChat [13], as can be seen in the supplementary material of the study by Ding et al. [5].
@@
Although most of these errors can easily be recognized and corrected by a pathologist, doing so adds to the workload again, undermining the potential benefits of automation. To address this problem, Ahmed et al. [1] applied a post-processing procedure using regular expressions based on a set of keywords to remove specific information, such as the precise anatomical location, from generated reports. However, this approach increases the system’s complexity for deployment, likely misses words that should be removed if the keyword set is not all-encompassing, and becomes more challenging to apply for removal of (sub)sentences.
@@
Furthermore, unlike most pathology-specific vision-language models, which as of yet have been trained primarily to generate diagnoses [1, 19], our focus lies on generating descriptions of cell and tissue patterns. The motivation for this is twofold: (1) writing these descriptions is often the most time-consuming part for a pathologist and could, for that reason, yield the largest efficiency gain if automated; and (2) reaching a definitive diagnosis for ambiguous cases can be difficult, if not impossible, without the results from additional diagnostic tests.
@@
As the main contribution of this work, we investigate how selecting information from pathology reports for vision-language modeling affects the quality of the multimodal representations and generated reports. Building upon a text preprocessing pipeline we developed in prior work [15], we compare training on full pathology reports against training only on the H&E-related sentences that describe the cell and tissue appearances. All experiments were performed using the BLIP-2 framework [9] and a dataset of H&E-stained WSIs with corresponding pathology reports for 19,636 cutaneous melanocytic lesions. Model performance was evaluated using image-to-text and text-to-image retrieval, as well as assessment of the accuracy and usability of the generated reports by an expert dermatopathologist. Moreover, all code and model parameters are made publicly available upon acceptance.
@@ 2.1 Dataset
The dataset used in this study consists of melanocytic lesion cases retrospectively collected from the digital archive of the Department of Pathology at the University Medical Center Utrecht, the Netherlands. All cases were accessioned between January 1, 2013, and December 31, 2020. More information about the curation process of the dataset can be found in Lucassen et al. [16]. For each case, all unique, H&E-stained WSIs and the corresponding pathology report were included after de-identification. The study was conducted in compliance with the hospital’s research ethics committee guidelines. Cases from patients who opted out of the use of their data for research purposes were excluded.
The pathology reports were preprocessed using a pipeline described in detail in prior work [15]. After translation from Dutch to English, the reports were segmented into subsentences based on the information content. Pretrained language models were used for the translation and segmentation after being finetuned for the respective tasks. Two variants of each report were created by selecting part of the subsentences: (1) all sentences from the original report; and (2) only the sentences with cell and tissue appearances written based on the H&E-stained WSIs. All cases with an empty report for one or both of the variants were excluded from the dataset.
Acquisition of the WSIs was performed using either a ScanScope XT scanner (Aperio, Vista, CA, USA) at 20$\times$ magnification with a resolution of 0.50 µm per pixel (slides scanned before 2016) or a NanoZoomer 2.0-XR scanner (Hamamatsu photonics, Hamamatsu, Shizuoka, Japan) at 40$\times$ magnification with a resolution of 0.23 µm per pixel (slides scanned starting from 2016). To guide the WSI tessellation, tissue cross-sections and pen markings were segmented in each WSI at 1.25$\times$ magnification using SlideSegmenter [14]. Non-overlapping tiles of 4,096$\times$4,096 pixels were extracted from the WSIs at 20$\times$ magnification. Tiles with identified pen markings or covered by tissue for less than 5% were excluded.
@@
Figure 1: Overview of the vision-language modeling framework. Feature vectors are extracted using HIPT [3] from all tiles of the tessellated WSIs. Pathology reports, with or without information selection, are tokenized and embedded. The Q-Former is trained for representation learning in the first stage and report generation in the second stage.
@@ 2.2 Vision-Language Model
We built upon the BLIP-2 framework [9], which was designed for parameter-efficient vision-language modeling. An overview of the vision-language model and training procedure is shown in Fig. 1.
All extracted WSI tiles for a case are converted to 192-dimensional feature vectors using the second stage of HIPT [3] (i.e., two successive Vision Transformers (ViTs) [6] pretrained on The Cancer Genome Atlas (TCGA) dataset [11]). This image encoder is connected using the so-called Querying Transformer (Q-Former) to a pretrained language model with mostly frozen parameters. The Q-Former consists of two parallel Transformer [21] submodules of 12 blocks: (1) an image submodule with trainable embeddings of 768 dimensions as input, also referred to as query embeddings, that extract information from the image feature vectors using cross-attention layers in every other block; and (2) a submodule with tokenized text embeddings as input and without cross-attention layers. Both submodules share the same self-attention layers, which enables interaction between image and text information through the query embeddings.
The output query embeddings returned by the image submodule of the Q-Former are prepended to the sequence of tokens for autoregressive language generation using BioGPT [17]. This language model has a decoder-only Transformer architecture with 24 layers, 347 million parameters, a vocabulary size of 42,384 tokens, and was pretrained on a biomedical text corpus.
@@ 2.3 Training for Representation Learning
To limit the required memory, only the Q-Former parameters were optimized for multimodal representation learning, while the parameters of the image encoder and text embedding layer remained frozen. The parameters of the Q-Former were initialized based on BERT-base-uncased [4], except for the cross-attention layers and query embeddings, which were randomly initialized at the start of training. Following the first stage of the BLIP-2 training procedure (see Fig. 1), the Q-Former was trained using an image-text contrastive loss $\mathcal{L}_{\text{ITC}}$, image-text matching loss $\mathcal{L}_{\text{ITM}}$, and image-grounded text generation loss $\mathcal{L}_{\text{ITG}}$.
@@
The image-grounded text generation loss is optimized by minimizing the cross-entropy of the paired text $\mathbf{y}$ under the forward autoregressive factorization with teacher-forcing [22] to parallelize computation:
@@
The model with 16 querying embeddings was trained on the sum of $\mathcal{L}_{\text{ITC}}$, $\mathcal{L}_{\text{ITM}}$, and $\mathcal{L}_{\text{ITG}}$ for 25 epochs with a batch size of 20. Model parameters were updated using a learning rate of 1 $\cdot$ 10 -4 with a cosine learning rate scheduler and 1,000 warmup steps. The AdamW [12] optimization algorithm ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) was used with weight decay equal to 0.01. The model was trained with label smoothing [20] for the contrastive loss with $\alpha$ = 0.9 (which was omitted from Eq. 1 for brevity). Hyperparameters were tuned based on the validation set results. The final model parameters were selected based on the epoch with the lowest validation loss.
@@ 2.4 Training for Report Generation
The model with 64 querying embeddings was trained for 21 epochs with a batch size of 36. Model parameters were updated using a learning rate of 1 $\cdot$ 10 -3 with a cosine learning rate scheduler and 1,000 warmup steps. The AdamW [12] optimization algorithm ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) was used with weight decay equal to 0.01. Hyperparameters were tuned based on the validation set results. The final model parameters were selected based on the epoch with the lowest validation loss.

### v3

@@ Abstract
Keywords: Vision Language Modeling, Histopathology, Text Preprocessing
@@ 1 Introduction
Vision-language modeling has seen much improvement in recent years (Li et al., 2023; Liu et al., 2024; Radford et al., 2021; Yu et al., 2022). Following the success in the domain of natural images, similar models have been developed for medical domains such as pathology (Ahmed et al., 2024; Ding et al., 2024; Lu et al., 2024; Shaikovski et al., 2024). Applications of vision-language models in pathology include uni- and cross-modal retrieval of cases from databases and automated report generation. The latter can potentially alleviate the increasing workload of pathologists (Berbís et al., 2023; Van der Laak et al., 2021).
@@
In addition to descriptions of cell and tissue appearances on hematoxylin and eosin (H&E)-stained whole slide images (WSIs), pathology reports often also include clinical information, patient history, and additional diagnostic results from immunohistochemical stains and molecular tests. These types of information are either difficult to predict correctly or cannot accurately be inferred from the H&E-stained WSIs at all. If sentences with this information are part of the training dataset, then vision-language models are prone to hallucination (i.e., the generation of statements that contradict or cannot be verified from the source content) (Ji et al., 2023). For example, outcomes of molecular testing are present in the reports generated by TITAN (Ding et al., 2024) in combination with PathChat (Lu et al., 2024).
@@
Although most of these errors can easily be recognized and corrected by a pathologist, doing so adds to the workload again, undermining the potential benefits of automation. To address this problem, Ahmed et al. (Ahmed et al., 2024) applied a post-processing procedure using regular expressions based on a set of keywords to remove specific information, such as the precise anatomical location, from generated reports. However, this approach increases the system’s complexity for deployment, likely misses words that should be removed if the keyword set is not all-encompassing, and becomes more challenging to apply for removal of (sub)sentences.
@@
Furthermore, unlike most pathology-specific vision-language models, which as of yet have been trained primarily to generate diagnoses (Ahmed et al., 2024; Shaikovski et al., 2024), our focus lies on generating descriptions of cell and tissue patterns. The motivation for this is twofold: (1) writing these descriptions is often the most time-consuming part for a pathologist and could, for that reason, yield the largest efficiency gain if automated; and (2) reaching a definitive diagnosis for ambiguous cases can be difficult, if not impossible, without the results from additional diagnostic tests.
@@
As the main contribution of this work, we investigate how selecting information from pathology reports for vision-language modeling affects the quality of the multimodal representations and generated reports. Building upon a text preprocessing pipeline we developed in prior work (Lucassen et al., 2024b), we compare training on full pathology reports against training only on the H&E-related sentences that describe the cell and tissue appearances. All experiments were performed using the BLIP-2 framework (Li et al., 2023) and a dataset of H&E-stained WSIs with corresponding pathology reports for 19,636 cutaneous melanocytic lesions. Model performance was evaluated using image-to-text and text-to-image retrieval, as well as assessment of the accuracy and usability of the generated reports by an expert dermatopathologist. Moreover, all code and model parameters are made publicly available  [^1] https://github.com/nuldertien/PathBLIP-2.
@@ 2.1 Dataset
The dataset used in this study consists of melanocytic lesion cases retrospectively collected from the digital archive of the Department of Pathology at the University Medical Center Utrecht, the Netherlands. All cases were accessioned between January 1, 2013, and December 31, 2020. More information about the curation process of the dataset can be found in (Lucassen et al., 2025). For each case, all unique, H&E-stained WSIs and the corresponding pathology report were included after de-identification. The study was conducted in compliance with the hospital’s research ethics committee guidelines. Cases from patients who opted out of the use of their data for research purposes were excluded.
The pathology reports were preprocessed using a pipeline described in detail in prior work (Lucassen et al., 2024b). After translation from Dutch to English, the reports were segmented into subsentences based on the information content. Pretrained language models were used for the translation and segmentation after being finetuned for the respective tasks. Two variants of each report were created by selecting part of the subsentences: (1) all sentences from the original report; and (2) only the sentences with cell and tissue appearances written based on the H&E-stained WSIs. All cases with an empty report for one or both of the variants were excluded from the dataset.
Acquisition of the WSIs was performed using either a ScanScope XT scanner (Aperio, Vista, CA, USA) at 20$\times$ magnification with a resolution of 0.50 µm per pixel (slides scanned before 2016) or a NanoZoomer 2.0-XR scanner (Hamamatsu photonics, Hamamatsu, Shizuoka, Japan) at 40$\times$ magnification with a resolution of 0.23 µm per pixel (slides scanned starting from 2016). To guide the WSI tessellation, tissue cross-sections and pen markings were segmented in each WSI at 1.25$\times$ magnification using SlideSegmenter (Lucassen et al., 2024a). Non-overlapping tiles of 4,096$\times$4,096 pixels were extracted from the WSIs at 20$\times$ magnification. Tiles with identified pen markings or covered by tissue for less than 5% were excluded.
@@
Figure 1: Overview of the vision-language modeling framework. Feature vectors are extracted using HIPT (Chen et al., 2022) from all tiles of the tessellated WSIs. Pathology reports, with or without information selection, are tokenized and embedded. The Q-Former is trained for representation learning in the first stage and report generation in the second stage.
@@ 2.2 Vision-Language Model
We built upon the BLIP-2 framework (Li et al., 2023), which was designed for parameter-efficient vision-language modeling. An overview of the vision-language model and training procedure is shown in Fig. 1.
All extracted WSI tiles for a case are converted to 192-dimensional feature vectors using the second stage of HIPT (Chen et al., 2022) (i.e., two successive Vision Transformers (ViTs) (Dosovitskiy et al., 2021) pretrained on The Cancer Genome Atlas (TCGA) dataset (Liu et al., 2018)). This image encoder is connected using the so-called Querying Transformer (Q-Former) to a pretrained language model with mostly frozen parameters. The Q-Former consists of two parallel Transformer (Vaswani et al., 2017) submodules of 12 blocks: (1) an image submodule with trainable embeddings of 768 dimensions as input, also referred to as query embeddings, that extract information from the image feature vectors using cross-attention layers in every other block; and (2) a submodule with tokenized text embeddings as input and without cross-attention layers. Both submodules share the same self-attention layers, which enables interaction between image and text information through the query embeddings.
The output query embeddings returned by the image submodule of the Q-Former are prepended to the sequence of tokens for autoregressive language generation using BioGPT (Luo et al., 2022). This language model has a decoder-only Transformer architecture with 24 layers, 347 million parameters, a vocabulary size of 42,384 tokens, and was pretrained on a biomedical text corpus.
@@ 2.3 Training for Representation Learning
To limit the required memory, only the Q-Former parameters were optimized for multimodal representation learning, while the parameters of the image encoder and text embedding layer remained frozen. The parameters of the Q-Former were initialized based on BERT-base-uncased (Devlin et al., 2019), except for the cross-attention layers and query embeddings, which were randomly initialized at the start of training. Following the first stage of the BLIP-2 training procedure (see Fig. 1), the Q-Former was trained using an image-text contrastive loss $\mathcal{L}_{\text{ITC}}$, image-text matching loss $\mathcal{L}_{\text{ITM}}$, and image-grounded text generation loss $\mathcal{L}_{\text{ITG}}$.
@@
The image-grounded text generation loss is optimized by minimizing the cross-entropy of the paired text $\mathbf{y}$ under the forward autoregressive factorization with teacher-forcing (Williams and Zipser, 1989) to parallelize computation:
@@
The model with 16 querying embeddings was trained on the sum of $\mathcal{L}_{\text{ITC}}$, $\mathcal{L}_{\text{ITM}}$, and $\mathcal{L}_{\text{ITG}}$ for 25 epochs with a batch size of 20. Model parameters were updated using a learning rate of 1 $\cdot$ 10 -4 with a cosine learning rate scheduler and 1,000 warmup steps. The AdamW (Loshchilov and Hutter, 2019) optimization algorithm ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) was used with weight decay equal to 0.01. The model was trained with label smoothing (Szegedy et al., 2016) for the contrastive loss with $\alpha$ = 0.9 (which was omitted from Eq. 1 for brevity). Hyperparameters were tuned based on the validation set results. The final model parameters were selected based on the epoch with the lowest validation loss.
@@ 2.4 Training for Report Generation
The model with 64 querying embeddings was trained for 21 epochs with a batch size of 36. Model parameters were updated using a learning rate of 1 $\cdot$ 10 -3 with a cosine learning rate scheduler and 1,000 warmup steps. The AdamW (Loshchilov and Hutter, 2019) optimization algorithm ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) was used with weight decay equal to 0.01. Hyperparameters were tuned based on the validation set results. The final model parameters were selected based on the epoch with the lowest validation loss.